# Identification and estimation of vaccine effectiveness in the test-negative design under equi-confounding

### v3

@@ 7 Sensitivity analysis
where $\eta$ is a scalar and $q(X)$ is a user-specified monotone sensitivity function. Together, they parameterize the deviation from equi-confounding on the odds ratio scale. This is an example of the well-known exponential tilt model [scharfstein_adjusting_1999, liu_identification_2020]. Setting $\eta=0$ corresponds to the case where equi-confounding holds, with values of $\eta$ further from zero representing larger violations. For example, one might choose $q(X)=1$, suggesting that confounding occurs uniformly across levels of $X$, and then vary $\eta$ over a grid of values in the interval $(-\omega,\omega)$.
@@ 8 Simulation
Likewise, violations of Assumption A2 are controlled by $\beta_{1V}$ and $\tau_{1V}$. When $\beta_{1V}\neq 0$ there is a direct effect of vaccination on test-negative illness $I=1$ and when $\tau_{1V}\neq 0$ there is a direct effect of vaccination on testing among those with test-negative illness. When this occurs, we can at most identify the ratio of vaccine effects on the two illness outcomes, i.e. $\exp(\beta_{2V}+\tau_{2V})/\exp(\beta_{1V}+\tau_{1V})$, provided there is no effect modification by $X$ or $U$. As we show in Appendix F and illustrate further via simulation below, under the special case of equal effects on testing, i.e. $\tau_{2V}=\tau_{1V}$, certain effects on symptomatic illness may still be recovered.
@@
Parameter values for all scenarios are shown in Table A1. For each scenario, we generate 1000 replicates and estimate $\Psi$ using the proposed estimators $\widehat{\Psi}_{om}^{*}$, $\widehat{\Psi}_{ipw}^{*}$, and $\widehat{\Psi}_{dr}^{*}$ as well as the conventional logistic regression estimator and calculate the bias and coverage of 95% confidence intervals based on the true value. For comparison, we also consider the following estimators based on data from the full cohort: a conventional outcome-modeling estimator adjusting for $X$ alone, another conditioning on $X$ and $U$ (covering the hypothetical case that all confounders are measured), and one based on $\Psi_{om}$ in Proposition 1. Finally, in scenario 8, we demonstrate the double robustness property of $\widehat{\Psi}_{dr}^{*}$ by adjusting the data generation process for $V$ and $I^{v}$ to include squared terms in $X$. Additional details on the setup and precise estimation methods are provided in Appendix I.
@@
The results are reported in Figures 2 and 3 with further details in Tables A3 and A4. When there is no unmeasured confounding, all estimators are unbiased. In all other scenarios the typical cohort estimator without $U$ is biased (scenario 1). When there is unmeasured confounding but our assumptions hold, the TND estimator is unbiased (scenario 2). When there is an effect of vaccination on the test-negative illness (scenario 3), the TND estimator is biased towards the null (when the effect is also protective); however, it does identify the ratio of vaccine effects against $I=2$ versus $I=1$. When equi-confounding or equi-selection are violated the TND estimator is biased (scenarios 4 and 5). When the test-negative and test-positive illnesses are not mutually exclusive, the TND estimator is biased for the causal risk ratio but unbiased for the causal odds ratio. When there is a direct effect of vaccination on testing behavior, the TND estimator remains unbiased if the effect is equivalent for test-negative and test-positive illnesses (scenario 6), but is biased when they are unequal (scenario 7). Finally, the proposed TND estimators, $\widehat{\Psi}_{om}^{*}$, $\widehat{\Psi}_{ipw}^{*}$, and $\widehat{\Psi}_{dr}^{*}$, remain unbiased in the presence of effect modification while the conventional logistic regression approach is biased for the marginal estimand (scenario 8). As expected, when the outcome model is misspecified $\widehat{\Psi}_{ipw}^{*}$ and $\widehat{\Psi}_{dr}^{*}$ remain unbiased while $\widehat{\Psi}_{om}^{*}$ is biased. Likewise, when the extended propensity score model is misspecified $\widehat{\Psi}_{om}^{*}$ and $\widehat{\Psi}_{dr}^{*}$ remain unbiased while $\widehat{\Psi}_{ipw}^{*}$ is biased. When both models are misspecified all estimators are biased.
@@ 9 Discussion
The plausibility of the equi-confounding assumption may be increased by knowledge of the source of the test-negative illness. In previous TNDs for influenza that included additional testing to confirm the source of test-negative illness the most common causes were rhinovirus, respiratory syncytial virus (RSV), human metapneumovirus (hMPV), parainfluenza virus (PIV), bocavirus (BoV), coronavirus (CoV), adenovirus (AdV), enterovirus (EV) and others [chua_use_2020-1]. Ideally, the test-negative illness should have a similar exposure mechanism and clinical presentation as the target illness but remain unaffected by the vaccine. For example, in pneumococcal disease, non-vaccine serotypes of Streptococcus pneumoniae serve as ideal test-negative controls [broome_pneumococcal_1980]. In Appendix H, we explore potential structural sources of equi-confounding —including health-seeking behavior and immortal time— as well as example violations, such as correlated vaccine-seeking [payne_impact_2023] and at home testing [qasmieh2024magnitude] and how these may be resolved.
@@
Even when equi-confounding does not hold, as our simulations suggest, the TND may still reduce bias relative to other approaches based only on a minimal covariate set, when test-negative illness serves as a reasonable, if imperfect, proxy for residual confounding. We also propose a sensitivity analysis that allows for straightforward departures from equi-confounding via the exponential tilt model. Alternatively, additional negative controls can be leveraged in a proximal inference framework to identify effects under more general unmeasured confounding structures [li_double_2023].
@@
Our results also depend on the assumption that the vaccine does not affect other infections or influence test-seeking behavior (except through infection). The former may be violated if vaccination provides short-term nonspecific protection through immune activation. Previous research has suggested that this is possible for influenza vaccines [cowling_increased_2012]. However, other results have shown the opposite [sundaram_influenza_2013]. Violations of this assumption would bias TND estimates, though in some cases, if the effect is in the same direction, the TND might still provide a lower bound on true vaccine effectiveness. If vaccination exerts equal effects on test-seeking behavior for test-positive and test-negative infections, as our simulations demonstrate, estimates of vaccine effectiveness against symptomatic infection remain unbiased.
[…]
$X$$V$$I=2$$T$$I=1$$U$A3bA3bA3aA3a
Figure 1: Causal directed acyclic graph (DAG) showing the assumed relationship between the variables in a test-negative study under equi-confounding. Here we split the $I$ node into $I=1$ and $I=2$ to show how the former acts as a negative outcome control. The equi-confounding assumption itself, as a parametric restriction, is not easy to represent on a DAG using standard notation. Here, we add color and labels over the arrows that must represent equal effects on the odds ratio scale for Assumptions A3a and A3b to hold.
Figure 2: Bias of vaccine effectiveness (VE) estimators across different simulation scenarios. In each scenario, we compare: (i) ideal cohort estimator with $U$ measured, (ii) more typical cohort estimator with $U$ unmeasured, (iii) a difference-in-differences (DiD) estimator in the cohort based on equation 6, and (iv) the TND estimator $\widehat{\Psi}_{om}^{*}$. Results are based on 1000 Monte Carlo replications. Black dashed line is the true VE. In scenario 3, grey dashed line is the ratio of VEs for test-positive versus test-negative illness. In scenarios 6 and 7, grey dashed line VE against symptomatic illness.
Figure 3: Bias of vaccine effectiveness (VE) estimators under effect heterogeneity and misspecification of the nuisance functions. We add to the estimators considered in Figure 2 the conventional TND logistic regression estimator (TND, logit) as well as the novel estimators of the marginal RRV introduced in section 6: the outcome modeling estimator (TND, om), the inverse probability weighting estimator (TND, ipw), and the doubly robust estimator (TND, dr). Results are based on 1000 Monte Carlo replications. Black dashed line is the true VE.
@@ Appendix A A nonparametric odds ratio model
[…]
@@
[…]
@@
Appendix B Proofs of main identifiability results
@@
    where $\text{where }OR_{y}(X)=\frac{\Pr(Y^{0}=y|V=1,X)\Pr(Y^{0}=0|V=0,X)}{\Pr(Y^{0}=0|V=1,X)\Pr(Y^{0}=y|V=0,X)}.$
@@ B.1 Proof of Proposition 1
$=\dfrac{\Pr(Y^{1}=2|V=1)}{E\left\{\Pr(Y^{0}=2|V=0,X)\dfrac{\Pr(Y^{0}=1|V=1,X)}{\Pr(Y^{0}=1|V=0,X)}\Big{|}V=1\right\}}$
$=\dfrac{\Pr(Y^{1}=2|V=1)}{E\left\{\Pr(Y^{0}=2|V=0,X)\dfrac{\Pr(Y^{1}=1|V=1,X)}{\Pr(Y^{0}=1|V=0,X)}\Big{|}V=1\right\}}$
$=\dfrac{\Pr(Y=2|V=1)}{E\left\{\Pr(Y=2|V=0,X)\dfrac{\Pr(Y=1|V=1,X)}{\Pr(Y=1|V=0,X)}\Big{|}V=1\right\}}$
@@ B.2 Proof of Proposition 2
Appendix C Additional Estimation Details
C.1 Cohort estimators
@@ C.3 Estimating Equations
[…]
@@
We can characterize the statistical properties of the estimator using M-estimation theory [stefanski_calculus_2002]. Under regularity conditions, we have the following asymptotic normality of the estimators:
[…]
@@
where $z_{\alpha}$ is the $100\alpha$th percentile of the standard normal distribution.
C.4 Bootstrapping
An alternative method for obtaining standard errors and confidence intervals for $\widehat{\Psi}^{*}_{om}$ or $\widehat{\Psi}^{*}_{ipw}$ is the bootstrap [diciccio_bootstrap_1987]. As in the main text, let $O=\{(X_{i},V_{i},S_{i}=1,Y_{i}^{*})\}_{i=1}^{n}$ represent the available data which are independent realizations from unknown probability distribution $F_{\theta}$.
@@ C.4 Bootstrapping
    where $z_{\alpha}$ is the $100\alpha$th percentile of the standard normal distribution.
@@
Appendix D Doubly robust estimator under cohort sampling
[…]

### v4

@@ 7 Sensitivity analysis
where $\eta$ is a scalar and $q(X)$ is a user-specified monotone sensitivity function. Together, they parameterize the deviation from equi-confounding on the odds ratio scale. This is an example of the well-known exponential tilt model [31, 32]. Setting $\eta=0$ corresponds to the case where equi-confounding holds, with values of $\eta$ further from zero representing larger violations. For example, one might choose $q(X)=1$, suggesting that confounding occurs uniformly across levels of $X$, and then vary $\eta$ over a grid of values in the interval $(-\omega,\omega)$.
@@ 8 Simulation
Likewise, violations of Assumption A2 are controlled by $\beta_{1V}$ and $\tau_{1V}$. When $\beta_{1V}\neq 0$ there is a direct effect of vaccination on test-negative illness $I=1$ and when $\tau_{1V}\neq 0$ there is a direct effect of vaccination on testing among those with test-negative illness. When this occurs, we can at most identify the ratio of vaccine effects on the two illness outcomes, i.e. $\exp(\beta_{2V}+\tau_{2V})/\exp(\beta_{1V}+\tau_{1V})$, provided there is no effect modification by $X$ or $U$. As we show in Appendix G and illustrate further via simulation below, under the special case of equal effects on testing, i.e. $\tau_{2V}=\tau_{1V}$, certain effects on symptomatic illness may still be recovered.
@@
Parameter values for all scenarios are shown in Table A1. For each scenario, we generate 1000 replicates and estimate $\Psi$ using the proposed estimators $\widehat{\Psi}_{om}^{*}$, $\widehat{\Psi}_{ipw}^{*}$, and $\widehat{\Psi}_{dr}^{*}$ as well as the conventional logistic regression estimator, $\widehat{\Psi}_{om}^{*}(X)$, and calculate the bias and coverage of 95% confidence intervals based on the true value. For comparison, we also consider the following estimators based on data from the full cohort: a conventional outcome-modeling estimator adjusting for $X$ alone, another conditioning on $X$ and $U$ (covering the hypothetical case that all confounders are measured), and one based on $\Psi_{om}$ in Proposition 1. Finally, in scenario 8, we demonstrate the double robustness property of $\widehat{\Psi}_{dr}^{*}$ by adjusting the data generation process for $V$ and $I^{v}$ to include squared terms in $X$. Additional details on the setup and precise estimation methods are provided in Appendix I.
@@
The results are reported in Figures 2 and 3 with further details in Tables A3 and A4. When there is no effect modification (scenarios 1 to 7) the proposed TND estimators and traditional logistic regression estimator are the same, so we show only the results from the latter, $\widehat{\Psi}_{om}^{*}(X)$, under the name “TND, logit”. When there is no unmeasured confounding, all estimators are unbiased (scenario 1). In all other scenarios, the cohort estimator without $U$ is biased. When there is unmeasured confounding but our assumptions hold, the TND estimator is unbiased (scenario 2). When there is an effect of vaccination on the test-negative illness (scenario 3), the TND estimator is biased towards the null (when the effect is also protective); however, it does identify the ratio of vaccine effects against $I=2$ versus $I=1$ (grey dashed line). When equi-confounding or equi-selection are violated the TND estimator is biased (scenarios 4 and 5). When there is a direct effect of vaccination on testing behavior, the TND estimator is unbiased for the risk ratio against symptomatic illness (grey dashed line) if the effect is equivalent for test-negative and test-positive illnesses (scenario 6, for more on this see Appendix G), but is biased when they are unequal (scenario 7). In all scenarios, the estimates from the TND estimator are the same as the difference-in-differences estimator (cohort, DiD) based on Proposition 1 applied to the full cohort even though the TND estimator uses only information among the tested, demonstrating how TND may be viewed as essentially a more efficient sampling design. Table A3 shows that, when unbiased, confidence intervals derived from estimating equations in Appendix D exhibit nominal coverage.
Finally, in the more realistic scenario, that the effect of vaccination is heterogeneous, the proposed TND estimators, $\widehat{\Psi}_{om}^{*}$$\widehat{\Psi}_{ipw}^{*}$, and $\widehat{\Psi}_{dr}^{*}$, remain unbiased for the marginal estimand while the conventional logistic regression approach, $\widehat{\Psi}_{om}^{*}(X)$, is biased (scenario 8). In Figure 3, we confirm the double robustness property of $\widehat{\Psi}_{dr}^{*}$. As expected, when the outcome model is misspecified $\widehat{\Psi}_{ipw}^{*}$ and $\widehat{\Psi}_{dr}^{*}$ remain unbiased while $\widehat{\Psi}_{om}^{*}$ is biased. Likewise, when the extended propensity score model is misspecified $\widehat{\Psi}_{om}^{*}$ and $\widehat{\Psi}_{dr}^{*}$ remain unbiased while $\widehat{\Psi}_{ipw}^{*}$ is biased. When both models are misspecified all estimators are biased.
@@ 9 Discussion
The plausibility of the equi-confounding assumption may be increased by knowledge of the source of the test-negative illness. In previous TNDs for influenza that included additional testing to confirm the source of test-negative illness the most common causes were rhinovirus, respiratory syncytial virus (RSV), human metapneumovirus (hMPV), parainfluenza virus (PIV), bocavirus (BoV), coronavirus (CoV), adenovirus (AdV), enterovirus (EV) and others [7]. Ideally, the test-negative illness should have a similar exposure mechanism and clinical presentation as the target illness but remain unaffected by the vaccine. For example, in pneumococcal disease, non-vaccine serotypes of Streptococcus pneumoniae serve as ideal test-negative controls [33]. However, more careful selection of test-negative infections based on source pathogen is currently limited by infrequent use of concurrent or multiplex testing. Care must also be taken when test-negative infections have associated vaccines as these can be sources of unequal confounding if not adjusted for.
@@
Even when equi-confounding does not hold, as our simulations suggest, the TND may still reduce bias relative to other approaches based only on a minimal covariate set, when test-negative illness serves as a reasonable, if imperfect, proxy for residual confounding. We also propose a sensitivity analysis that allows for straightforward departures from equi-confounding via the exponential tilt model. Alternatively, additional negative controls can be leveraged in a proximal inference framework to identify effects under more general unmeasured confounding structures [34].
@@
Our results also depend on the assumption that the vaccine does not affect other infections or influence test-seeking behavior (except through infection). The former may be violated if vaccination provides short-term nonspecific protection through immune activation. Previous research has suggested that this is possible for influenza vaccines [35]. However, other results have shown the opposite [36]. Violations of this assumption would bias TND estimates, though in some cases, if the effect is in the same direction, the TND might still provide a lower bound on true vaccine effectiveness. If vaccination exerts equal effects on test-seeking behavior for test-positive and test-negative infections, as our simulations demonstrate, estimates of vaccine effectiveness against symptomatic infection remain unbiased.
Lastly, our results have implications for other study designs for vaccine evaluation. First, [37] recently proposed a cluster-randomized TND that corrects for potential bias caused by intervention-induced differential health-seeking behavior in an unblinded trial. In our framework, this corresponds to a setting where Assumption A3a holds by design, as vaccination is randomized, but Assumption A3b does not, as individuals are aware of their assignment and therefore may differentially seek care when sick. As mentioned above, to remove the influence of intervention effects on health-seeking requires them be equal across test-positive and test-negative illnesses on the multiplicative scale. Second, in settings where full cohort data are available, comparing test-positive and test-negative infections via Expressions 6 or 7 offers an alternative approach to confounding control, allowing for estimation of absolute risks under different vaccination regimes.
@@ Appendix A A nonparametric odds ratio model
Following [20], we parameterize a unit’s contribution to the likelihood for the potential outcome under no treatment $Y^{0}$ conditional on $V$ and covariates $X$ and assuming independent sampling. Let
@@
$f(Y^{0}=y|V=1,X=x)=\dfrac{h(y,x)\exp\{\beta(y,x)\}}{\int h(y,x)\exp\{\beta(y,x)\}\;dy}$
where $\int h(y,x)\exp\{\beta(y,x)\}\;dy$ is a proportionality or normalizing constant, ensuring the left-hand side is a proper density. This parameterization can, in principle, represent any likelihood function one might encounter in practice and therefore imposes no restrictions on the data generation process (i.e. it is completely nonparameteric).
@@
Appendix B Extended discussion of equi-confounding assumption
B.1 Unmeasured health-seeking behavior
Among the stated motivations for the test-negative design when it was first introduced was the potential for the TND to adjust for unmeasured health-seeking behavior. Early observational studies of flu vaccines tended to overstate their effectiveness relative to randomized trials. At the time, it was hypothesized that the bias may have been caused by differential health-seeking behavior, as healthier adults were more likely to get vaccinated and also engage in other health behaviors, such as hand-washing and personal distancing, that reduce risk of infection. While hard to measure as a traditional covariate, similar health-seeking might also be revealed by an individual’s propensity to pursue a test when sick and therefore conditioning on receipt of a test may effectively resolve the issue.
Under our framework, if unmeasured health-seeking exerts equal influence on the risk of infection for test-positive and test-negative illness then it satisfies Assumption A3a, i.e.
$\dfrac{\Pr(I^{0}=2|V=1,X)}{\Pr(I^{0}=2|V=0,X)}=\dfrac{\Pr(I^{0}=1|V=1,X)}{\Pr(I^{0}=1|V=0,X)}.$
This may be more plausible when the other risk-reducing interventions that these individuals engage in are not specific to the focal illness, but instead affect risk of other infections similarly.
At the same time, if unmeasured health-seeking behavior results in equal propensity for pursuing a test when sick then it also satisfies Assumption A3b
$\dfrac{\Pr(T^{0}=1|I^{0}=2,V=1,X)}{\Pr(T^{0}=1|I^{0}=2,V=0,X)}=\dfrac{\Pr(T^{0}=1|I^{0}=1,V=1,X)}{\Pr(T^{0}=1|I^{0}=1,V=0,X)}.$
The plausibility of this assumption can be increased by certain design elements of the TND. Namely, the symptom screen, when well-defined and effectively implemented, equalizes the conditions under which a test is pursued while the (assumed) unavailability of testing outside health facilities ensures individuals are blind to the causative agent.
We note that the equi-confounding framework also allows for alternative equi-confounding structures involving unmeasured health-seeking that could lead to over- or under-optimistic estimates of vaccine effectiveness.
B.2 Immortal time among vaccinated
Concerns have also been raised about the potential for design-induced biases in the test-negative design resulting from failure to explicitly emulate a randomized trial [29]. Here we show that, when the degree of design-induced bias is similar for test-positive and test-negative illnesses, they tend to cancel each other out and produce no bias in the TND estimate of vaccine effectiveness when they are equivalent.
$\frac{\Pr(Y^{0}=2|V=1,X)}{\Pr(Y^{0}=2|V=0,X)}=\delta_{2}(X)$
for some $\delta_{2}(X)<1$ as the person time in the first 28 days after receiving a vaccine is differentially discarded among those who received a vaccine compared to those who did not. However, if this exclusion of person time among the vaccinated is applied independently of the test result, the bias also applies to the incidence of test-negative illness such that
$\frac{\Pr(Y^{0}=1|V=1,X)}{\Pr(Y^{0}=1|V=0,X)}=\delta_{1}(X)$
for some $\delta_{1}(X)<1$. In the special case that $\delta_{2}(X)=\delta_{1}(X)$, which could occur if the trends in the incidence of test-positive and test-negative illness among the unvaccinated are parallel over follow up (before and after the 28 day period where immortal time is accruing), we have equi-confounding and the design-induced bias is completely removed by $\Psi^{*}_{om}$ and $\Psi^{*}_{ipw}$.
B.3 Correlated vaccination history
Prior work on the TND suggests that the design could be biased when there are other vaccines covering test-negative illness in addition to the focal vaccine for the test positive pathogen when vaccination behavior is correlated. For instance, denote by $V^{*}$ the vaccination against the test-negative pathogen and assume first $V^{*}$ is unavailable to the investigator. As shown in the DAG in Figure A1 this would also violate our identifiability assumptions because it would imply the existence of a confounder that does not affect the test positive and test negative illnesses equally, i.e.,
B.4 Home testing
Alternatively, multiplex rapid diagnostics tests that test for multiple pathogens simultaneously are increasingly available for home use. If one could limit test-negative controls to illnesses covered by multiplex testing then rapid at-home testing may affect test-seeking equally for test-positive and test-negative illnesses (unless one is perceived as more dangerous/life-threatening), in which case it may no longer bias our estimates.
[…]
(a) Correlated vaccination behavior
XXVVI=2I=2RRTTI=1I=1UU
(b) At home testing
Appendix C Proofs of main identifiability results
@@
    $\text{where }OR_{y}(X)=\frac{\Pr(Y^{0}=y|V=1,X)\Pr(Y^{0}=0|V=0,X)}{\Pr(Y^{0}=0|V=1,X)\Pr(Y^{0}=y|V=0,X)}.$
@@ C.1 Proof of Proposition 1
$=\dfrac{\Pr(Y^{1}=2|V=1)}{E\left\{\int\mathbbm{1}(Y^{0}=2)\dfrac{f(Y^{0}=y|V=0,X)\exp\{\beta(y,X)\}}{\int f(Y^{0}=y|V=0,X)\exp\{\beta(y,X)\}dy}dy\Big{|}V=1\right\}}$
@@
[…]
@@ C.2 Proof of Proposition 2
Appendix D Additional Estimation Details
D.1 Cohort estimators
@@ D.3 Estimating Equations
We first introduce additional notation. As before, we observe data $O=(X,V,S=1,Y^{*})$ for $i=1,\ldots,n$, where $n$ is the number of observed units. Let $\mathbb{P}(g)$ denote the average of function $g(O)$ across $n$ units, i.e., $\mathbb{P}(g)=n^{-1}\sum_{i=1}^{n}g(O_{i})$. Let $\operatorname{expit}(x)=\exp(x)/{1+\exp(x)}$. For random variables $V$ and $W$ , let $V\overset{P}{\rightarrow}W$ and $V\overset{D}{\rightarrow}W$ denote $V$ converges to $W$ in probability and in distribution, respectively.
@@
We can characterize the statistical properties of the estimator using M-estimation theory [38]. Under regularity conditions, we have the following asymptotic normality of the estimators:
$\sqrt{n}\left\{\big{(}\widehat{\Psi}^{*},\widehat{\nu}\big{)}^{\prime}-\big{(}\underline{\Psi}^{*},\underline{\nu}\big{)}^{\prime}\right\}\xrightarrow{D}N\left(0,V_{1}^{-1}V_{2}\left(V_{1}^{-1}\right)^{\prime}\right)$
@@
[…]
D.4 Bootstrapping
An alternative method for obtaining standard errors and confidence intervals for $\widehat{\Psi}^{*}_{om}$ or $\widehat{\Psi}^{*}_{ipw}$ is the bootstrap [39]. As in the main text, let $O=\{(X_{i},V_{i},S_{i}=1,Y_{i}^{*})\}_{i=1}^{n}$ represent the available data which are independent realizations from unknown probability distribution $F_{\theta}$.
@@ D.4 Bootstrapping
[…]
@@
Appendix E Doubly robust estimator under cohort sampling
Recall, the parameter of interest is the causal risk ratio among the vaccinated $\Psi=\Pr(Y^{1}=2|V=1)/\Pr(Y^{0}=2|V=1)$. We focus on the denominator $\psi=\Pr(Y^{0}=2|V=1)$ as identification of the numerator is trivial under causal consistency. Identification of $\psi$ under equi-confounding is structurally similar to Universal Difference-in-Differences (UDiD), with a key difference that the NCO and Outcome are mutually exclusive by design. Under the UDiD framework, [20] provide the following estimators of $\psi$: